OpenStudy (anonymous):

Find the derivative of f(x) = negative 7 divided by x at x = -3

OpenStudy (decentnabeel):

\[\frac{d}{dx}\left(\frac{-7}{x}\right)=\frac{7}{x^2}\] then x replace -3 \[\frac{7}{\left(-3\right)^2}=\frac{7}{9}\]

OpenStudy (anonymous):

Thanks

OpenStudy (decentnabeel):

no problem @Nairz

OpenStudy (jacob902):

example Find the derivative of f(x) = negative 9 divided by x at x = -8. write it in index form f(x)=−9x−1 now find the derivative lastly substitute x = -8 into the derivative

OpenStudy (jacob902):

example Find the derivative of f(x) = negative 9 divided by x at x = -8. write it in index form f(x)=−9x−1 now find the derivative lastly substitute x = -8 into the derivative